OpenStudy (anonymous):

The sum of the squares of two consecutive positive even integers is one hundred. Find the two integers.

OpenStudy (mertsj):

x^2+(x+2)^ 2= 100 x^2+x^2+4x+4=100 2x^2+4x-96=0 x^2+2x-48=0 (x+8)(x-6)=0 x=-8 0 or 6 The problem specifies positive so the required integers are 6 and 8

OpenStudy (anonymous):

try some numbers and see if it works

OpenStudy (anonymous):

satellite can u come help me

OpenStudy (anonymous):

for example 4 and 6 don't work because \[4^2+6^2=16+36\] so try 6 and 8

OpenStudy (anonymous):

We usually use these numbers in right triangle for hypotenuse. 3^2+4^2=5^2 6^2+8^2=10^2 5^2+12^2=13^2